Draw all possible samples of size 3 with replacement from a population consisting of 3, 6, 9, 12, and 15 and calculate the sample means.


1
Expert's answer
2021-07-16T01:49:22-0400

N=5

n=3

Number of samples "= N^n = 5^3 = 125"